I attempted this to solve this expression and came up with log(a)(y+2x+7). Is this correct. If not please help solve.
Write the expression log(a)(y+5)+2log(a)(x+1) as one logarithm.

Answer by venugopalramana(3286)

I attempted this to solve this expression and came up with log(a)(y+2x+7). Is this correct. If not please help solve.
Write the expression log(a)(y+5)+2log(a)(x+1)
USE LOG (X*Y)=LOG(X)+LOG(Y)...AND....LOG(X^N)=N*LOG(X)
log(a)(y+5)+2log(a)(x+1) =LOG(a)(Y+5)+LOG(a)(X+1)^2=LOG(a){(Y+5)(X+1)^2} IS THE ANSWER.
as one logarithm.